OpenStudy (anonymous):

Please factor the following trinomial. w^18 – 9w^9y^5 + 14y^10

OpenStudy (anonymous):

factor is with GCF right?

OpenStudy (anonymous):

yes I do believe

OpenStudy (anonymous):

is there a space between – 9w^9y^5?

OpenStudy (anonymous):

so it would be -9w^9 Y^5

OpenStudy (anonymous):

\[w ^{18} - 9w ^{9}y ^{5} + 14y ^{10}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[w^{18} - 9w^9y^5 + 14y^{10}\] So, let: \[w^9 = x\] So expression becomes: \[x^2 - 9xy^5 + 14y^{10}\] Now can you do factorization??

OpenStudy (anonymous):

i think that you multiply the 2 outer #'s to get 14

OpenStudy (anonymous):

and 7 times 2 is 14 and adds up to 9

OpenStudy (anonymous):

Yes so these will be the factors... -7 and -2

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

I don't know where to begin with this.. I'm sorry. I just don't like this kind of math

OpenStudy (anonymous):

\[x^2 - 7y^5.x - 2y^5.x + 14y^{10} = x(x - 7y^5) - 2y^5(x - 7y^5) = (x-2y^5)(x-7y^5)\] Now put you know x = w^9 Put there to get the answer..

OpenStudy (anonymous):

Ya hes right

OpenStudy (anonymous):

thank you guys very much!

OpenStudy (anonymous):

your very welcome

OpenStudy (anonymous):

Welcome dear..